OpenStudy (anonymous):

P(blue or white) blue is 10 white is 3 write each Probability as a fraction in simplest form

OpenStudy (anonymous):

p(b) = 10/13 p(w) = 3/13 add together 13/13 i'm not sure i understood the problem though

OpenStudy (anonymous):

ok so this is the question. Suppose you collect data on the color of shirts worn by the students in your class. Eight wore green,five wore red, three wore white, and ten wore blue. Write each probability as a fraction in simplest form. altogether the socks count as 26

OpenStudy (anonymous):

P(green)= #of that color shirt (in this case green)/total shirts= 8/26 (you can simplify this further since both 8 and 26 are even numbers so you can divide both by 2). P(red)= same idea as green P(white)= same idea again P(blue)= same type of thing as above So when you find out what the P(blue) and P(white) is you can add their fractions together just like how pi.i said.